Two positive charges, each of 4.8 µC, and negative charge -6.36 µC are fixed at the vertices of an equilateral triangle of side 13 cm. Find the electrical force on negative charge.
1
Expert's answer
2020-11-06T10:14:04-0500


We are given

"q_1=q_2=q=+4.8 \cdot 10^{-6}C"

"q_3=-6.36 \cdot 10^{-6}C"

"r=0.13m"

Force acting on charge q3 from the side of charge q1

"F_{13}=\frac{1}{4 \pi\epsilon_0 } \cdot \frac{q_1 \cdot q_3}{r^2}=\frac{1}{4 \pi\epsilon_0 } \cdot \frac{q \cdot q_3}{r^2}"

Force acting on charge q3 from the side of charge q2

"F_{23}=\frac{1}{4 \pi\epsilon_0 } \cdot \frac{q_2 \cdot q_3}{r^2}=\frac{1}{4 \pi\epsilon_0 } \cdot \frac{q \cdot q_3}{r^2}"

"F_{13}=F_{23}=F_0"

Then behind the cosine theorem we write

"F^2=F_{13}^2+F_{23}^2-2 \cdot F_{13} \cdot F_{23} \cdot \cos(\gamma)"

"F^2=F_{13}^2+F_{23}^2-2 \cdot F_{13} \cdot F_{23} \cdot \cos(180^0-2 \alpha)"

"F^2=2F_{0}^2+2 F_{0}^2 \cdot \cos(2 \alpha)=2F_0^2 \cdot (1+\cos(60^0))"

"F^2=2F_0^2 \cdot (1+\frac{1}{2})=2F_0^2 \cdot \frac{3}{2}=3 \cdot F_0^2"

"F=\sqrt {3} \cdot F_0=\sqrt {3} \cdot\frac{1}{4 \pi\epsilon_0 } \cdot \frac{q \cdot q_3}{r^2}"


"F=\sqrt {3} \cdot\frac{1}{4 \pi\cdot 8.854 \cdot 10^{-12} } \cdot \frac{4.8 \cdot 10^{-6} \cdot 6.36 \cdot 10^{-6}}{0.13^2}=28.12N"